OpenStudy (sedatefrog712):

MEDAL A Labrador leaps over a hurdle. The function f(t) represents the height of the Labrador above the ground, in inches, at t seconds: f(t) = -16t2 + 20t A foxhound jumps over the same hurdle. The table shows the height of the foxhound above the ground g(t), in inches, at t seconds: Time (t) g(t) 0 0 0.4 7.04 0.6 8.64 0.75 9 1.0 8 1.5 0

OpenStudy (sedatefrog712):

Part A: Compare and interpret the maximum of f(t) and g(t)? (4 points) Part B: Which function has a greater x-intercept? What do the x-intercepts of the graphs of f(t) and g(t) represent? (4 points) Part C: Determine the y-intercepts of both functions, and explain what this means in the context of the problem. (2 points)

OpenStudy (sedatefrog712):

@One098

OpenStudy (sedatefrog712):

i need help

OpenStudy (sedatefrog712):

not really but thanx anyway

OpenStudy (sedatefrog712):

@ganeshie8

OpenStudy (sedatefrog712):

can you help me please

ganeshie8 (ganeshie8):

which part ?

OpenStudy (sedatefrog712):

all of them

ganeshie8 (ganeshie8):

Part A: Compare and interpret the maximum of f(t) and g(t)? (4 points)

ganeshie8 (ganeshie8):

Look at the table of foxhound, whats its max height ?

OpenStudy (sedatefrog712):

8.64

ganeshie8 (ganeshie8):

Time (t) g(t) 0 0 0.4 7.04 0.6 8.64 0.75 `9` 1.0 8 1.5 0

ganeshie8 (ganeshie8):

did you miss 9 ?

OpenStudy (sedatefrog712):

yeah

ganeshie8 (ganeshie8):

so the maximum height reached by foxhound/g(t) is `9`

ganeshie8 (ganeshie8):

what about Labrador/f(t) ?

ganeshie8 (ganeshie8):

\(\large f(t) = -16t^2 + 20t\)

OpenStudy (sedatefrog712):

when i graph it it say 6.25

ganeshie8 (ganeshie8):

do you know how to find the maximum height of this parabola ?

ganeshie8 (ganeshie8):

You're right! so the labrador/f(t) reaches only a max height of 6.25

OpenStudy (sedatefrog712):

not really im not good at parabola

OpenStudy (sedatefrog712):

but i guess im good at graphing

ganeshie8 (ganeshie8):

thats very good :) so are we done with part a

OpenStudy (sedatefrog712):

ok can you help me with part b

OpenStudy (sedatefrog712):

how would i find the x-intercepts on this?

ganeshie8 (ganeshie8):

if you want to find the max height reached by labrador/f(t) algenraically : \(\large f(t) = -16t^2 + 20t\) \(\large -b/2a = -20/-2*16 = 5/8\) max height = \(\large f(5/8) = -16(5/8)^2 + 20(5/8) = 25/4 = 6.25\)

ganeshie8 (ganeshie8):

that was still for part a ^^

OpenStudy (sedatefrog712):

ok that makes sense

ganeshie8 (ganeshie8):

So clearly the foxhound reached greatest height.. thats the interpretation ok

ganeshie8 (ganeshie8):

lets see part b

ganeshie8 (ganeshie8):

Part B: Which function has a greater x-intercept? What do the x-intercepts of the graphs of f(t) and g(t) represent? (4 points)

ganeshie8 (ganeshie8):

x-intercept occurs when the function becomes 0

ganeshie8 (ganeshie8):

Time (t) g(t) `0 0` 0.4 7.04 0.6 8.64 0.75 9 1.0 8 `1.5 0 `

ganeshie8 (ganeshie8):

the x intercepts of foxhound/g(t) are 0, 1.5 they tell you that the foxhound has started leaping at t=0 and fell at t=1.5

ganeshie8 (ganeshie8):

so the foxhound is in air for 1.5 seconds. is that clear ?

ganeshie8 (ganeshie8):

can you find the x intercepts for labrador/f(t) ? \(\large f(t) = -16t^2 + 20t\)

OpenStudy (sedatefrog712):

sorry lag my internet

ganeshie8 (ganeshie8):

its okie

ganeshie8 (ganeshie8):

just set that expression equal to 0 and solve t

ganeshie8 (ganeshie8):

\( \large -16t^2 + 20t = 0\)

OpenStudy (sedatefrog712):

so would i replace t for 0 into f(t)

ganeshie8 (ganeshie8):

factor the GCF and solve t

ganeshie8 (ganeshie8):

not yet. we will be doing that for part c, for y intercepts.

ganeshie8 (ganeshie8):

for x intercepts, you need to solve above equation

OpenStudy (sedatefrog712):

the gcf is 4 correct

ganeshie8 (ganeshie8):

4t

ganeshie8 (ganeshie8):

or may be -4t

OpenStudy (sedatefrog712):

176

ganeshie8 (ganeshie8):

\(\large -16t^2 + 20t = 0 \) \(\large -4t(4t -5) = 0 \) \(\large t = 0\) or \(\large t = 5/4 = 1.25\)

ganeshie8 (ganeshie8):

so the labrador is in air for exactly 1.25 seconds

OpenStudy (sedatefrog712):

oh ok im sorry i got a little confused man im an idiot

ganeshie8 (ganeshie8):

no you're not

ganeshie8 (ganeshie8):

it seems foxhound is in air for longer time than the labrador ?

ganeshie8 (ganeshie8):

np :) can you finish part c on ur own ?

ganeshie8 (ganeshie8):

Part C: Determine the y-intercepts of both functions, and explain what this means in the context of the problem. (2 points)

ganeshie8 (ganeshie8):

y intercept the value of function at t=0

ganeshie8 (ganeshie8):

Time (t) g(t) `0 0` 0.4 7.04 0.6 8.64 0.75 9 1.0 8 1.5 0

OpenStudy (sedatefrog712):

actually i need help on that i jumped the gun a bit

ganeshie8 (ganeshie8):

for foxhound/g(t) the x intercept is 0, that means the foxhound started from ground, from 0 height

OpenStudy (sedatefrog712):

oh ok and its the same thing for the Labrador correct

ganeshie8 (ganeshie8):

what about the y intercept of labrador/f(t) ?

ganeshie8 (ganeshie8):

yup! but u need to show why its the same

OpenStudy (sedatefrog712):

f(t)=-16(0)^2 + 20(0)

ganeshie8 (ganeshie8):

Perfect!

ganeshie8 (ganeshie8):

so they both have started from the ground at 0 height

OpenStudy (sedatefrog712):

Yay thank you so much you are the best

ganeshie8 (ganeshie8):

yw:)